what is the value of expressions?
15-3+2-9=?
16+4:2+9=?

Answer by stanbon(48546)

what is the value of expressions?
15-3+2-9=?
17-12 = ?
5 = ?
==============
16+4:2+9=?
Comment: Your ":" confuses the problem.
Cheers,
Stan H.